OpenStudy (anonymous):

DETERMINING THE FORMULA FOR HYDRATES HELP!! Cocl2 ? H20 mass of hydrate : anhydrous mass : mass of water : moles cocl2 moles h20 ratio: formula name : I really don't get this .. can someone explain

OpenStudy (cuanchi):

I assume you have an experimental value for the mass of Hydrate and for the anhydrous mass. The difference between this two mass will be the mass of water CoCl2 ? H20 y (58.9 + 2 (35.5)) + x (2 +16) = mass of hydrate x = moles of H2O = mass of water/ 18 y = moles of CoCl2 = mass of anhydrous/129.9 Ratio = x/y formula = y CoCl2 x H2O ( the most probable formula is "CoCl2 6 H2O" , but depends of your experimental data this values for x and y can be different.) name Cobalt(II) chloride hexahydrate if you are interested more about this compound you can consult wikipedia http://en.wikipedia.org/wiki/Cobalt(II)_chloride

OpenStudy (anonymous):

thank you !